OpenStudy (anonymous):

Which measure of central tendency best describes this situation: the favorite fruit sold in the cafeteria? Explain. Which measure of central tendency best describes this situation: the favorite fruit sold in the cafeteria? Explain. @Mathematics

OpenStudy (anonymous):

Median; there likely are no outliers. Mean; there likely are no outliers. Median; there likely are outliers. Mode; the data are not numerical. ^ These are the answers!

OpenStudy (wasiqss):

median i guess because median gives value with the the highest frequency

OpenStudy (wasiqss):

n the fruit that is sold with highest frequency will ultimately be the fav one

OpenStudy (anonymous):

So you think it is the first one or the third one? <3

OpenStudy (wasiqss):

jus temme the meanin of outliers then i ll tell ,i forgot the meaning

OpenStudy (anonymous):

You tell me [: I am lost on this whole thing! Google doesnt know either! / :

OpenStudy (wasiqss):

ahhh i told u the concept now lemme google it

OpenStudy (anonymous):

11. You buy a shirt that costs $9.95. With tax, you pay $10.65. Write and solve an equation to find the amount of tax x you paid. (2 points) 9.95 – x = 10.65; $0.70 9.95 + 10.65 = x; $20.60 9.95 + x = 10.65; $0.70 10.65 + x = 9.95; $0.70

OpenStudy (anonymous):

Forget that one I will figure it out, can you help me with this one?

OpenStudy (wasiqss):

even google is unable to answer ur ques :P

OpenStudy (anonymous):

Your telling me. Because my stuff is to smart for Google. I am kidding! xD

OpenStudy (wasiqss):

the third one is rite in the equation because 9.95 is net amount n u add tax of x to get 10.65

OpenStudy (wasiqss):

n yehh i wish i was smarter than google to solve the above ques :D

OpenStudy (anonymous):

12. the length of a swimming pool 20 km 20 m 20 cm 20 mm

OpenStudy (wasiqss):

20m, i know cox i swim daily :D

OpenStudy (anonymous):

[: I have a dentist app, and I have like 50 more problems / : This is what I need help with so just try your best!

OpenStudy (anonymous):

Thankyou [:

OpenStudy (anonymous):

8. The formula F = + 37 estimates the temperature F in degrees Fahrenheit when crickets chirp n times per minute. Use the formula to estimate the temperature when n = 75 chirps/min. Round to the nearest whole number, if necessary. 58°F 53°F 52°F 56°F

OpenStudy (anonymous):

For the fruit question, it's the mode because you need numbers to calculate everything else and types of fruit are not numbers... :)

OpenStudy (anonymous):

F=4/n+37 *

OpenStudy (anonymous):

6. Find the outlier in the data: 32, 40, 34, 31, 20, 36, 40, 31. How does the outlier affect the mean? (2 points) 45; raises it by 1.7. 20; lowers it by 1.9. 45; lowers it by 1.7. 20; raises it by 1.9.

OpenStudy (anonymous):

5. The number of patients treated at Dr. McCormick's dentist office each day was recorded for nine days. These are the data: 20, 9, 8, 17, 19, 16, 13, 14, 8. Find the mean, median, and mode of the data. If necessary, round to the nearest tenth. (2 points) 14, 11.8, 8 11.8, 14, 8 13.8, 8, 14 13.8, 14, 8

OpenStudy (anonymous):

I can probally figure out the rest if you guys can help me with these last 2. Thankyou Jemurray3

OpenStudy (wasiqss):

andrea cool down baby :P

OpenStudy (anonymous):

Im trying, I still have to get ready & everything. Gr.

OpenStudy (anonymous):

C:

OpenStudy (anonymous):

Folding clothes, Doing math...... Dentist in an hour to get wisdom teeth cut out!!!! NO FUN!

OpenStudy (wasiqss):

add all the 9 digits and divide by 9 to get mean mode is the number which occurs the most among 9 number and for median arrange all the numb in sequence and the fifth numb will be median as it is middle value

OpenStudy (wasiqss):

awwww andrea poor u :P

OpenStudy (anonymous):

I know! I will wait till my Dad gets home to help me so I can call him [:

OpenStudy (wasiqss):

n yeah listen for ques 6 the outlier is 20, outlier means that data is outta range or sequence, so to get the difference include 20 in ur mean, and then take out mean without including 20, so the diff btw 2 mean will lead u to answer

OpenStudy (wasiqss):

n yeh u in which grade?

OpenStudy (wasiqss):

u on fb?

OpenStudy (anonymous):

10th.

OpenStudy (anonymous):

So it would be raised or lowered. And yes I have a facebook?

OpenStudy (wasiqss):

tell ur facebook name, as u can always message n send me ques, n yeh it wud be lowered because the out of range num is lower than sequence

OpenStudy (anonymous):

I dont give out my facebook!

OpenStudy (anonymous):

13. 22.62 mm = __ m (2 points) 0.2262 0.02262 0.002262 22,620

OpenStudy (anonymous):

15. 0.155 mL = __ L [:

OpenStudy (wasiqss):

awww y??